# Morphology-Specific Peptide Discovery via Masked Conditional Generative Modeling

### v2

@@ Results
All peptides selected for simulation-based validation were run in triplicate (90 simulations in total). In every run, the peptides formed aggregates: each exceeded the aggregation-propensity threshold ($\mathrm{AP}\geq 1.8$, computed from the CG-MD SASA ratio) and showed visible assembly upon trajectory inspection. With respect to morphology targeting, the RMOI distributions show distinct tendencies for the two classes: spherical assemblies cluster around high values ($\sim 0.8$), while fibrillar assemblies concentrate at low values ($\sim 0.3$), as seen in Figure 3d,e. This separation highlights that RMOI captures the expected contrast between compact and elongated morphologies, albeit only as a proxy. Nevertheless, the cutoff values were empirically defined, and the metric has significant limitations. We observed several borderline cases where RMOI classified assemblies incorrectly, despite visual inspection confirming the target morphology. Because RMOI considers only the longest and shortest inertia directions, it neglects how mass is distributed in the simulation box — for instance, aggregates aligned along box boundaries in all directions sometimes yielded high RMOI despite being fiber-like. More broadly, RMOI struggles to reliably classify fibrils: slightly wider fibers tend to exhibit intermediate values, and the metric inherently overemphasizes tubular/rod geometries rather than fibrillar ones. By visually inspecting the final frame of each trajectory (see Supplementary Material), we determine a morphology success rate of 83%.
@@
Compared with validation under training-like conditions using our Gaussian Mixture Model procedure, enforcing the target morphology descriptors leads to a substantially larger drop in valid generated peptides. This suggests that the chosen condition combinations are both underrepresented in the training distribution and intrinsically difficult, yielding low data support and fewer sequences that satisfy all constraints. The effect is particularly pronounced for fibrillar targets: the required $\beta$-sheet fraction occurs in $0.1\%$ of training peptides (Figure 1f), which further depresses the hit rate.
@@
Figure 4: Latent map of morphology conditioning and morphology matching results. A UMAP is fit on the conditional prior centers with encoder posteriors of generated peptides projected into the same space (dots for spheres, stars for fibers). (a) Condition center embeddings are colored by conditioned sequence length, revealing an ordered length gradient along the sphere branch and a single, compact, and distinct fiber island; notably, fiber targets of length 7 cluster closer to the sphere branch, consistent with over-constrained fiber conditions claims. (b) Same embedding with MD-simulated peptides highlighted according to target morphology and outcome (success/failure under the RMOI criterion); stars mark the sphere/fiber prior centers. (c) Morphology success rates for the 30 candidates under the RMOI rule and independent visual assessment were reported by target morphology. In total, RMOI success is based on the mean across three runs, with visual success on the majority vote.
@@
This visualization contextualizes the morphology-specific conditioning: fiber-targeted conditions collapse into a compact, sparsely populated island, consistent with the rarity of $\beta$-sheet-positive peptides, whereas sphere-targeted conditions span several sub-modes, reflecting broader but structured support. Even so, it is important to recall that the validated sphere set remains compositionally narrow, with amino acid usage dominated by F/I/L/V (Figure 3b), suggesting that the descriptor choices together with the AP $\geq\!1.8$ and structural filters bias toward a restricted set of chemistries. A closer look at the fiber samples reveals a connection between the sequence length of seven in the region corresponding to spheres and the island corresponding to true fiber conditions. This suggests that conditioning for shorter sequences with $\beta$-sheet content, which was used as a target for fibers, is likely an overconstrained requirement (possibly unrealistic), as the resulting sequences do not cluster near their intended condition centers. This reinforces both the challenges of overconstraining and the importance of masking and range exploration in maintaining valid generative support.
Discussion
@@
We have developed PepMorph, an end-to-end framework for generating novel, diverse, aggregation-prone peptides, explicitly conditioned on geometric and physicochemical descriptors that govern their assembly morphology. Our pipeline is extremely successful at generating highly novel and diverse sequences, with $83\%$ of them adopting the intended morphology under targeted conditioning. PepMorph addresses a gap in peptide generative modeling by enabling user-specified morphology control through a transformer-based CVAE with a masking mechanism for flexible conditioning on peptide descriptors, whereas prior peptide generators have not explicitly modeled self-assembly [13, 14] or provided any morphology goal or conditioning [21].
@@
A key advantage of PepMorph is that its conditioning mechanism is both descriptor-agnostic and mask-aware: any peptide descriptor with an assignable value (and, optionally, a predictor for validation) can be integrated without changing the core architecture. This makes the framework naturally extensible along two axes. First, it can easily scale in the sequence space: extending to longer peptides primarily requires additional data and model capacity, while the masked conditioning and autoregressive decoding remain directly applicable. Second, it can expand in the morphology space: the same interface can accommodate alternative morphologies by using different and possibly richer descriptors, enabling targets beyond spheres and fibrils (e.g., sheets or nets). The main challenge here lies in quantitatively characterizing more complex morphologies — an inherently difficult task, as illustrated by the limitations of RMOI. Beyond morphology, application-driven properties such as antimicrobial activity, previously used in generative peptide design [32, 14], can also be seamlessly incorporated. In short, PepMorph provides a reusable scaffold for expanding from short peptides and two morphologies to broader sequence lengths and richer structural or functional targets.
While the current descriptor set works well for the chosen targeted conditioning, two practical issues are worth improving. First, PEP-FOLD occasionally fails to produce conformations for specific sequences, leading to missing 3D-derived descriptors and, consequently, discarded candidates. Further adjustments on PEP-FOLD or even entirely replacing it with more robust structure predictors would reduce the amount of discarded candidates. Second, some regions of descriptor space are considerably underrepresented (e.g., sequences with any $\beta$-sheet content are $\sim 0.1\%$, with almost all of them having 10 residues — see Supplementary Material), which can bias learning and make those targets harder to satisfy. Curating additional data in these sparse regimes and performing expert-guided descriptor selection are straightforward, high-yield next steps that complement the existing pipeline.
@@
Regarding the model, the simple approach of modeling the conditional prior as a single diagonal Gaussian works well in our setting, as reflected by strong condition matching performance in validation. However, under partial conditioning with a sparse mask $m$, the mapping from descriptors to valid sequences is inherently multimodal. A unimodal prior averages across disparate modes, which weakens constraint satisfaction and can cause the decoder to default to frequent training patterns. A promising direction for future work would be to develop architectures that adopt a richer masked prior, such as explicitly multimodal or mixture-based formulations [33, 34], making the uncertainty induced by masked descriptors explicit, rather than collapsing it into a single Gaussian.
@@
Since the aggregation labels in our dataset were obtained with CG-MD, we validated candidate morphologies in silico, using a similar CG-MD setup, rather than experiments. This choice ensures methodological consistency, but MD predictions are known to be sensitive to the employed force field and simulation setting, and do not always match the experimental predictions. Thus, re-labeling the aggregation labels with more accurate molecular models would be very impactful. Particularly promising in this respect are machine-learning interatomic potentials [35, 36] — especially implicit-solvent variants [37, 38, 39] — which can approach ab-initio accuracy while retaining near-linear scaling with system size. Integrating such models would yield higher-fidelity training and validation labels while keeping the end-to-end design cycle computationally tractable. In addition, the simulation–generation loop could be tightened by calculating and conditioning directly on morphology descriptors such as RMOI.
@@ Methods
All peptide–level 3D descriptors are computed from the lowest-energy structure predicted by PEP-FOLD [24]. The code was executed in parallel and returns the top five models per peptide; it then selects the model with the best internal score in PDB format. Failed predictions were retried up to three times. From each selected PDB, we parsed backbone and side-chain coordinates using Biopython’s PDBParser [40]. Residue secondary structure is assigned with the DSSP algorithm [41]. Peptides for which DSSP or PDB coordinate parsing failed were omitted from conditioning and, consequently, from model training/validation. All descriptor–extraction steps were parallelized. We compute the $\beta$-sheet fraction as
@@
where $L_{E}$ is the number of residues labeled E (extended strand) by DSSP. Because nonzero $f_{\beta}$ values are rare in our dataset, we binarized this feature as a flag indicating $\beta$-sheet content if $f_{\beta}>0$.
For the hydrophobic moment, we use the Eisenberg hydrophobicity scale $h_{l}$ for residue $l$ [42] and define $\hat{\mathbf{v}}_{l}$ as the vector along $C_{\alpha}\to C_{\beta}$ for residue $l$ (for Gly or missing $C_{\beta}$, a unit vector is used). The hydrophobic-moment vector and its magnitude are
@@
We fine-tune a lightweight, two-head predictor on top of an ESM-2 encoder (t12/35M) [29]. Given an input peptide $y=(y_{1},\dots,y_{L})$, with $L_{\max}=10$, where each $y_{l}$ is an amino acid symbol, we tokenize the input for ESM-2 and take the final ESM layer representation, mask out special and padding tokens, and compute a masked mean sequence representation. A linear projection maps this representation to a shared hidden state, from which two heads produce (i) a scalar aggregation propensity $\hat{a}$ and (ii) a probability $\hat{p}$ for assembly. The objective loss of the model is a sum of a binary cross-entropy term on $\hat{p}$ and a mean-squared error term on $\hat{a}$. We do training in two stages: (i) firstly, we freeze the ESM-2 encoder and train only the heads; (ii) then, we unfreeze and fine-tune end-to-end with discriminative learning rates. As such, we firstly train the heads with AdamW (lr $10^{-3}$) for 5 epochs; we then unfreeze the ESM-2 encoder and fine-tune with discriminative learning rates (encoder $10^{-5}$, heads $10^{-3}$) and an exponential learning rate decay for 6 more epochs.
@@
with $\beta$ a cyclic Kullback–Leibler (KL) weight and $\lambda_{\text{mask}},\lambda_{\text{bin}},\lambda_{\text{cont}}$ scalar hyperparameters. The reconstruction loss $\mathcal{L}_{\mathrm{rec}}$ is the negative log–likelihood of the ground–truth amino–acid at each non-<pad> position under the decoder’s categorical distribution (token–level cross–entropy with label smoothing and teacher forcing).
For the divergence term, we follow the VAEAC approach [15] and use the closed‐form KL divergence between diagonal Gaussians. Let the encoder posterior for sample $n$ be $q_{n}(z\mid y)=\mathcal{N}(\mu_{n},\operatorname{diag}(\sigma_{n}^{2}))$ and the masked conditional prior be $p_{n}(z\mid s)=\mathcal{N}(\mu_{\mathrm{prior},n},\operatorname{diag}(\sigma_{\mathrm{prior},n}^{2}))$, both in latent dimension $K$. The KL loss is then
@@
i.e., the same binary cross–entropy, but applied only to the observed binary descriptors for each sample. For the continuous descriptors, let $\mathcal{C}_{n}=\{i\in\{1,\dots,d\}:m_{n,i}=1\}$ be the continuous descriptors observed for sample $n$, with target value $c_{n,i}$ and prediction $\hat{c}_{n,i}$. We define
@@
To maximize the performance of the model and better attain the main goals of generation — broad exploration of sequence space and strong compliance with arbitrary partial conditions —, we introduce three adjustments to the training data usage. Firstly, we attenuate class imbalance in the binary descriptors with weighted sampling ($\times 2$ for positive is_assembled, $\times 10$ for positive has_beta_sheet_content). Adding to this, in order to expose the model to partial conditions, we apply stochastic masking during training: for each example, a random subset of currently available descriptors is set to unobserved ($m_{i}=0$), ensuring at least one descriptor remains observed. Finally, each epoch is augmented with $5,000$ uniformly sampled short peptides (up to $L_{\max}$) that observe only the length descriptor, seeking to achieve robustness at encoding the entirety of the peptide space.
@@
Individual PDB files were built from the FASTA sequences via Avogadro v1.2 [43] using geometry optimization. All CG-MD simulations used GROMACS [44] with MARTINI 3 [45] forcefield, applying bead substitutions (Q5 to Q4, TC5 to SC4) following Sasselli and Coluzza’s adaptation to small peptides [46]. Topologies were generated with martinize2 [47] from the all‐atom PDBs to obtain the corresponding coarse-grained representations.
Each system comprised $300$ peptides in a $15\times 15\times 15~\mathrm{nm}^{3}$ cubic, periodic box [23]. The boxes were then solvated with the MARTINI water model and neutralized by adding Na+/Cl- when needed to achieve electroneutrality. Energy minimization employed the steepest‐descent integrator for $5,000$ steps, with reaction‐field electrostatics and $1.2$ nm cutoffs for both Coulomb and Van der Waals (VdW) interactions. Production simulations consisted of 20 million steps ($\Delta t$ = 25 fs, totalizing 500 ns) in the NPT ensemble at 303 K and 1 bar. A frame checkpoint of the trajectory is stored every 50k steps (1.25 ns). We used the velocity‐rescale thermostat, C-rescale barostat and Particle–Mesh Ewald for electrostatics with a cutoff of $1.2$ nm. VdW interactions used a 1.2 nm cutoff with the potential-shift-Verlet modifier. Neighbor search used the Verlet scheme with a 0.005 buffer tolerance. Each system was run 3 times with different initial velocities for statistical purposes. All simulation runs used GPU acceleration.

### v3

@@ 2 Results
All 30 peptides yielded by PepMorph for simulation-based validation, as well as additional control cohorts totalizing 87 peptides, were simulated in triplicate under the same CG-MD protocol and analysis pipeline. For PepMorph-selected peptides, aggregation was observed in every run: each exceeded the aggregation-propensity threshold ($\mathrm{AP}\geq 1.8$, computed from the CG-MD SASA ratio) and showed visible assembly upon trajectory inspection. With respect to morphology targeting, the RMOI distributions show distinct tendencies for the two classes: spherical assemblies cluster around high values ($\sim 0.8$), while fibrillar assemblies concentrate at low values ($\sim 0.3$), as seen in Figure 3d,e. This separation highlights that RMOI captures the expected contrast between compact and elongated morphologies, albeit only as a proxy. Nevertheless, the cutoff values were empirically defined, and the metric has significant limitations. We observed several borderline cases where RMOI classified assemblies incorrectly, despite visual inspection confirming the target morphology. Because RMOI considers only the longest and shortest inertia directions, it neglects how mass is distributed in the simulation box — for instance, aggregates aligned along box boundaries in all directions sometimes yielded high RMOI despite being fiber-like. More broadly, RMOI struggles to reliably classify fibrils: slightly wider fibers tend to exhibit intermediate values, and the metric inherently overemphasizes tubular/rod geometries rather than fibrillar ones. By visually inspecting the final frame of each trajectory (see Supplementary Material), we determine a morphology success rate under our CG-MD validation protocol of 83% (Table 6). Visual labeling used explicit geometric criteria: spherical denotes a single compact aggregate with low apparent anisotropy, whereas fibrillar denotes an elongated aggregate with a dominant long axis; ambiguous or sheet-/net-like assemblies were labeled as failures for the targeted class.
As previously mentioned, and to contextualize this CG-MD-based validation and provide negative baselines, we additionally selected multiple control cohorts. These isolate (i) aggregation enrichment from the AP/SA screening stage and (ii) morphology steering from the proxy-descriptor windows: we selected random peptides (n=15) and non-morphology-proxy (unconditional) generated peptides (n=12; two per length), which provide lower-bound baselines with no design intent; we then selected cohorts that failed the AP filter (AP-fail), which test whether aggregation is obtained without the aggregation screen (n=30; n=15 per morphology); and then cohorts that passed the AP filter but failed the descriptor match, which test whether aggregation alone is sufficient to recover the intended morphologies without morphology-proxy constraints (n=30; n=15 per morphology).
As summarized in Table 6, AP-fail cohorts aggregate rarely, while descriptor-fail cohorts often aggregate but achieve the intended morphologies at markedly lower rates than PepMorph candidates, indicating that aggregation enrichment and morphology steering are separable effects in the pipeline. To quantify the strength and uncertainty of morphology steering, we compare PepMorph candidates against the aggregation-matched descriptor-fail controls using Wilson 95% confidence intervals for success rates and two-sided Fisher exact tests, reported in Supplementary Material. By visual criteria, PepMorph exhibits substantially higher morphology success, with odds ratios of $2.5$ ($p=0.410$) for spheres and $7.6$ ($p=0.042$) for fibrils. If we pool across the two targeted tasks, it yields an odds ratio of $4.3$ ($p=0.022$). These results support that the proxy-descriptor windows enrich for the intended morphologies beyond aggregation screening alone, while the modest MD cohort sizes imply wide uncertainty, particularly for the spherical target.
Table 6: CG-MD aggregation and morphology outcomes for PepMorph candidates and control cohorts. Aggregation is defined by AP $\geq 1.8$ (majority across three runs). For targeted cohorts, morphology success rates (RMOI and visual) are reported conditional on aggregation. RMOI success uses a majority vote across three runs given thresholds (spheres $\geq 0.75$, fibrils $\leq 0.35$), and visual success is also based on majority vote across three runs. For untargeted cohorts (random/unconditional), no morphology target is defined; we therefore report, among aggregating sequences, the fraction whose aggregates fall into sphere-like (S) or fibril/tube-like (F) regimes.
Cohort Aggregation rate RMOI outcome Visual outcome
Untargeted
Random (n=15) 20.0% (3/15) S: 0.0% (0/3) F: 0.0% (0/3) S: 0.0% (0/3) F: 0.0% (0/3)
Unconditional (n=12) 58.3% (7/12) S: 14.3% (1/7) F: 57.1% (4/7) S: 14.3% (1/7) F: 71.4% (5/7)
Spherical-targeted
Failed AP screen (n=15) 6.7% (1/15) 0.0% (0/1) 0.0% (0/1)
Failed descriptor screen (n=15) 86.7% (13/15) 46.2% (6/13) 61.5% (8/13)
PepMorph candidates (n=15) 100.0% (15/15) 73.3% (11/15) 80.0% (12/15)
Fibril-targeted
Failed AP screen (n=15) 26.7% (4/15) 0.0% (0/4) 50.0% (2/4)
Failed descriptor screen (n=15) 86.7% (13/15) 23.1% (3/13) 46.2% (6/13)
PepMorph candidates (n=15) 100.0% (15/15) 26.7% (4/15) 86.7% (13/15)
PepMorph (all) (n=30) 100.0% (30/30) 50.0% (15/30) 83.3% (25/30)
We also performed targeted robustness controls, fully displayed and contextualized in Supplementary Material, on a randomly selected subset of validated peptides (8 per target; triplicate) to probe sensitivity to to force-field choice in aggregation propensity (MARTINI 2.2 versus MARTINI 3), and periodic-boundary artifacts for morphology (larger box at matched concentration). Aggregation propensity remains in the same regime across setups, supporting that aggregation propensity is comparatively transferable across MARTINI 2.2 and MARTINI 3. In contrast, morphology is more sensitive: spherical outcomes remain largely consistent under the larger-box condition, whereas fibrillar outcomes show some setup sensitivity, yielding networked or amorphous aggregates. Some of these results could still, however, be attributed to short and yet-to-converge trajectories, especially for the larger box, which emphasizes the extensive computational cost associated with larger setups. Even so, this motivates interpreting the MD-based morphology validation as protocol-conditional and highlights that proxy selection and validation protocol jointly determine end-to-end outcomes.
@@
As it pertains to the overall throughput of the pipeline, compared with validation under training-like conditions using our Gaussian Mixture Model procedure, enforcing the target morphology descriptors leads to a substantially larger drop in valid generated peptides. This suggests that the chosen condition combinations are both underrepresented in the training distribution and intrinsically difficult, yielding low data support and fewer sequences that satisfy all constraints. The effect is particularly pronounced for fibrillar targets: the required $\beta$-strand assignmet occurs in $0.1\%$ of training peptides (Figure 1f), which further depresses the hit rate.
@@
Figure 4: Latent map of morphology conditioning. A UMAP is fit on the conditional prior centers with encoder posteriors of generated peptides projected into the same space (dots for spheres, stars for fibers). (a) Condition center embeddings are colored by conditioned sequence length, revealing an ordered length gradient along the sphere branch and a single, compact, and distinct fiber island; notably, fiber targets of length 7 cluster closer to the sphere branch, consistent with over-constrained fiber conditions claims. (b) Same embedding with MD-simulated peptides highlighted according to target morphology and outcome (success/failure under the RMOI criterion); stars mark the sphere/fiber prior centers.
@@
This visualization contextualizes the morphology-specific conditioning: fiber-targeted conditions collapse into a compact, sparsely populated island, consistent with the rarity of $\beta$-strand-positive peptides, whereas sphere-targeted conditions span several sub-modes, reflecting broader but structured support. Even so, it is important to recall that the validated sphere set remains compositionally narrow, with amino acid usage dominated by hydrophobic residues like F/I/L/V (Figure 3b), suggesting that the descriptor choices together with the AP $\geq\!1.8$ and structural filters bias toward a restricted set of chemistries. In particular, an atomistic follow-up on two highly hydrophobic sphere-target sequences, present in Supplementary Material Section Targeted fully atomistic re-simulations, results in fibrillar aggregation, while all CG-MD runs — including with the MARTINI 2.2 and larger box runs — yielded a spherical aggregate. This illustrates that high hydrophobic moment alone does not prevent fibrillar outcomes when overall hydrophobicity is extreme, motivating future proxy refinement. A closer look at the fiber samples reveals a connection between the sequence length of seven in the region corresponding to spheres and the island corresponding to true fiber conditions. This suggests that conditioning for shorter sequences with $\beta$-strand assignment, which was used as a target for fibers, is likely an overconstrained requirement (possibly unrealistic), as the resulting sequences do not cluster near their intended condition centers. This reinforces both the challenges of overconstraining and the importance of masking and range exploration in maintaining valid generative support.
3 Discussion
@@
We have developed PepMorph, an end-to-end framework for generating novel, diverse, aggregation-prone peptides, explicitly conditioned on geometric and physicochemical descriptors that inform their assembly morphology. Our pipeline is extremely successful at generating highly novel and diverse sequences, with $83\%$ of them adopting the intended morphology under targeted conditioning. PepMorph addresses a gap in peptide generative modeling by enabling user-specified proxy-guided morphology control through a transformer-based CVAE with a masking mechanism for flexible conditioning on peptide descriptors, whereas prior peptide generative models have predominantly focused on functional activities such as antimicrobial properties dasPepCVAESemiSupervisedTargeted2018, szymczakDiscoveringHighlyPotent2023, and existing self-assembly-oriented generators have not provided any morphology steering as part of the design objective njirjakReshapingDiscoverySelfassembling2024a.
@@
A key advantage of PepMorph is that its conditioning mechanism is both descriptor-agnostic and mask-aware: any peptide descriptor with an assignable value (and, optionally, a predictor for validation) can be integrated without changing the core architecture. This makes the framework naturally extensible along two axes. First, it can easily scale in the sequence space: extending to longer peptides primarily requires additional data and model capacity, while the masked conditioning and autoregressive decoding remain directly applicable. Second, it can expand in the morphology space: the same interface can accommodate alternative morphologies by using different and possibly richer or higher-fidelity descriptors, enabling targets beyond spheres and fibrils (e.g., sheets or nets). The main challenge here lies in quantitatively characterizing more complex morphologies — an inherently difficult task, as illustrated by the limitations of RMOI. Beyond morphology, application-driven properties such as antimicrobial activity, previously used in generative peptide design liFoundationModelIdentifies2024, szymczakDiscoveringHighlyPotent2023, can also be seamlessly incorporated. In short, PepMorph provides a reusable scaffold for expanding from short peptides and two morphologies to broader sequence lengths and richer structural or functional targets.
While the current descriptor set demonstrates morphology steering under our chosen targeting window, two practical issues are worth improving. First, the choice of proxy descriptors, as well as their feasible target ranges, materially affect end-to-end success. In the present work, morphology control is therefore best interpreted as protocol-conditional: it is validated under a fixed CG-MD setup and morphology criterion, and additional sensitivity checks indicate that outcomes can shift under protocol changes (e.g., larger boxes) or higher-resolution atomistic settings for specific sequences. This highlights that expert-guided refinement of the proxy set — for example, adding complementary constraints such as an explicit hydrophobicity cap for sphere targeting — may improve robustness without altering the core partial-conditioning framework. Second, some regions of descriptor space are considerably underrepresented (e.g., sequences with any $\beta$-strand assignment are $\sim 0.1\%$, with almost all of them having 10 residues — see Supplementary Material), which can bias learning and make those targets harder to satisfy. Curating additional data in these sparse regimes and performing expert-guided descriptor selection are straightforward, high-yield next steps that complement the existing pipeline.
@@
Regarding the model, the simple approach of modeling the conditional prior as a single diagonal Gaussian works well in our setting, as reflected by strong condition matching performance in validation. However, under partial conditioning with a sparse mask $m$, the mapping from descriptors to valid sequences is inherently multimodal. A unimodal prior averages across disparate modes, which weakens constraint satisfaction and can cause the decoder to default to frequent training patterns. A promising direction for future work would be to develop architectures that adopt a richer masked prior, such as explicitly multimodal, mixture-based or flow-based prior formulations wangDiverseAccurateImage2017, lavdaImprovingVAEGenerations2019, leeDesignPeptidesNoncanonical2025, making the uncertainty induced by masked descriptors explicit, rather than collapsing it into a single Gaussian.
@@
Since the aggregation labels in our dataset were obtained with CG-MD, we validated candidate morphologies in silico, using a similar CG-MD setup, rather than experiments. This choice ensures methodological consistency, but MD predictions are known to be sensitive to the employed force field and simulation setting, and do not always match the experimental predictions. Thus, re-labeling the aggregation labels with more accurate molecular models would be very impactful. Particularly promising in this respect are machine-learning interatomic potentials thalerLearningNeuralNetwork2021, rockenAccurateMachineLearning2024 — especially implicit-solvent variants rockenPredictingSolvationFree2025, costeDevelopingImplicitSolvation2024, thalerDeepCoarsegrainedPotentials2022 — which can approach ab-initio accuracy while retaining near-linear scaling with system size. Integrating such models would yield higher-fidelity training and validation labels while keeping the end-to-end design cycle computationally tractable. In addition, the simulation-generation loop could be tightened by calculating and conditioning directly on morphology descriptors such as RMOI. Ultimately, more extensive atomistic simulations or even experimental characterization would be required to establish morphology as ground truth and to calibrate which proxy descriptors and simulation protocols transfer reliably across conditions. This would be a natural next step to extend and further validate this work.
@@ 4 Methods
All peptide–level 3D descriptors are computed from the lowest-energy structure predicted by PEP-FOLD reyPEPFOLD4PHdependentForce2023. The code was executed in parallel and returns the top five models per peptide; it then selects the model with the best internal score in PDB format. Failed predictions were retried up to three times. From each selected PDB, we parsed backbone and side-chain coordinates using Biopython’s PDBParser cockBiopythonFreelyAvailable2009. Residue secondary structure is assigned with the DSSP algorithm hekkelmanMkdsspCalculateSecondary. Peptides for which DSSP or PDB coordinate parsing failed were omitted from conditioning and, consequently, from model training/validation — unless AP or SA/no-SA labels were defined; in that case, the masking mechanism handles the missing proxy descriptors. All descriptor–extraction steps were parallelized. We compute the monomeric $\beta$-strand assignment as
@@
where $L_{E}$ is the number of residues labeled E (extended strand) by DSSP. Because nonzero $f_{\beta}$ values are rare in our dataset, we binarized this feature as a flag indicating $\beta$-strand content if $f_{\beta}>0$.
For the hydrophobic moment, we use the Eisenberg hydrophobicity scale $h_{l}$ for residue $l$ eisenbergAnalysisMembraneSurface1984 and define $\hat{\mathbf{v}}_{l}$ as the vector along $C_{\alpha}\to C_{\beta}$ for residue $l$ (for Gly or missing $C_{\beta}$, a unit vector is used). The hydrophobic-moment vector and its magnitude are
@@
We fine-tune a lightweight, two-head predictor on top of an ESM-2 encoder (t12/35M) linLanguageModelsProtein2022a. Given an input peptide $y=(y_{1},\dots,y_{L})$, with $L_{\max}=10$, where each $y_{l}$ is an amino acid symbol, we tokenize the input for ESM-2 and take the final ESM layer representation, mask out special and padding tokens, and compute a masked mean sequence representation. A linear projection maps this representation to a shared hidden state, from which two heads produce (i) a scalar aggregation propensity $\hat{a}$ and (ii) a probability $\hat{p}$ for assembly. The objective loss of the model is a sum of a binary cross-entropy term on $\hat{p}$ and a mean-squared error term on $\hat{a}$. We do training in two stages: (i) firstly, we freeze the ESM-2 encoder and train only the heads; (ii) then, we unfreeze and fine-tune end-to-end with discriminative learning rates. As such, we firstly train the heads with AdamW (lr $10^{-3}$) for 5 epochs; we then unfreeze the ESM-2 encoder and fine-tune with discriminative learning rates (encoder $10^{-5}$, heads $10^{-3}$) and an exponential learning rate decay for 6 more epochs.
@@
with $\beta$ a cyclic Kullback-Leibler (KL) weight and $\lambda_{\text{mask}},\lambda_{\text{bin}},\lambda_{\text{cont}}$ scalar hyperparameters. The reconstruction loss $\mathcal{L}_{\mathrm{rec}}$ is the negative log-likelihood of the ground-truth amino-acid at each non-<pad> position under the decoder’s categorical distribution (token-level cross-entropy with label smoothing and teacher forcing).
For the divergence term, we follow the VAEAC approach ivanovVariationalAutoencoderArbitrary2018 and use the closed‐form KL divergence between diagonal Gaussians. Let the encoder posterior for sample $n$ be $q_{n}(z\mid y)=\mathcal{N}(\mu_{n},\operatorname{diag}(\sigma_{n}^{2}))$ and the masked conditional prior be $p_{n}(z\mid s)=\mathcal{N}(\mu_{\mathrm{prior},n},\operatorname{diag}(\sigma_{\mathrm{prior},n}^{2}))$, both in latent dimension $K$. The KL loss is then
@@
i.e., the same binary cross-entropy, but applied only to the observed binary descriptors for each sample. For the continuous descriptors, let $\mathcal{C}_{n}=\{i\in\{1,\dots,d\}:m_{n,i}=1\}$ be the continuous descriptors observed for sample $n$, with target value $c_{n,i}$ and prediction $\hat{c}_{n,i}$. We define
@@
To maximize the performance of the model and better attain the main goals of generation — broad exploration of sequence space and strong compliance with arbitrary partial conditions —, we introduce three adjustments to the training data usage. Firstly, we attenuate class imbalance in the binary descriptors with weighted sampling ($\times 2$ for positive is_assembled, $\times 10$ for positive has_beta_strand). Adding to this, in order to expose the model to partial conditions, we apply stochastic masking during training: for each example, a random subset of currently available descriptors is set to unobserved ($m_{i}=0$), ensuring at least one descriptor remains observed. Finally, each epoch is augmented with $5,000$ uniformly sampled short peptides (up to $L_{\max}$) that observe only the length descriptor, seeking to achieve robustness at encoding the entirety of the peptide space.
@@
Conditional Likelihood Evaluation
To quantify conditional generalization of the decoder, we compute teacher-forced token-level negative log-likelihood (NLL) on a held-out test set under different conditioning masks. For each test peptide $y=(y_{1},\dots,y_{L})$ with available descriptor vector $c$ and mask $m$, we form the condition summary $s=\phi([c\odot m,\,m])$ and run the decoder in teacher-forcing mode, i.e., at step $t$ the model is given the ground-truth prefix $(y_{1},\dots,y_{t-1})$ and predicts a categorical distribution over $y_{t}$. We compute the per-token cross-entropy (in nats) over non-<pad> tokens as
$\mathrm{CE}(y;c,m)=-\frac{1}{L^{\prime}}\sum_{t\in\mathcal{T}}\log p_{\theta}\!\big(y_{t}\mid y_{<t},\,s\big),$ (11)
where $\mathcal{T}$ indexes all non-<pad> target positions (including <eos>) and $L^{\prime}=|\mathcal{T}|$. We report the mean CE across the test set, and its corresponding perplexity $\mathrm{PPL}=\exp(\mathrm{CE})$.
We evaluate three mask settings: (i) observed-mask, using each test sample’s native availability mask $m$ (reflecting naturally missing descriptors); (ii) full-mask on the complete-case subset (dataset subset where all descriptors are observed, $m=\mathbf{1}$); and (iii) random-on-top on the same complete-case subset, where we additionally drop each descriptor independently with probability $0.5$ (while ensuring at least one descriptor remains observed) and compute CE under the resulting mask.
@@
Individual PDB files were built from the FASTA sequences via Avogadro v1.2 hanwellAvogadroAdvancedSemantic2012 using geometry optimization. All CG-MD simulations used GROMACS abrahamGROMACSHighPerformance2015 with MARTINI 3 souzaMartini3General2021 forcefield (except the 16 simulations for AP control verification, that leveraged MARTINI 2.2 talluriDiscoveryUnconventionalNonintuitive2025), applying bead substitutions (Q5 to Q4, TC5 to SC4) following Sasselli and Coluzza’s adaptation to small peptides sasselliAssessmentMARTINI32024. Topologies were generated with martinize2 kroonMartinize2VermouthUnified2025 from the all‐atom PDBs to obtain the corresponding coarse-grained representations.
Each system comprised $300$ peptides in a $15\times 15\times 15~\mathrm{nm}^{3}$ cubic, periodic box wangAggregationRulesShort2024, except the 16 simulations for box-size effects, which utilized a $24\times 24\times 24~\mathrm{nm}^{3}$ box, with the same concentration ($1230$ peptides). The boxes were then solvated with the MARTINI water model and neutralized by adding Na+/Cl- when needed to achieve electroneutrality. Energy minimization employed the steepest‐descent integrator for $5,000$ steps, with reaction‐field electrostatics and $1.2$ nm cutoffs for both Coulomb and Van der Waals (VdW) interactions. Production simulations consisted of 20 million steps ($\Delta t$ = 25 fs, totalizing 500 ns) in the NPT ensemble at 303 K and 1 bar. A frame checkpoint of the trajectory is stored every 50k steps (1.25 ns). We used the velocity‐rescale thermostat, C-rescale barostat and Particle–Mesh Ewald for electrostatics with a cutoff of $1.2$ nm. VdW interactions used a 1.2 nm cutoff with the potential-shift-Verlet modifier. Neighbor search used the Verlet scheme with a 0.005 buffer tolerance. Each system was run 3 times with different initial velocities for statistical purposes. All simulation runs used GPU acceleration.
All-Atom Molecular Dynamics Simulations
We performed two targeted AA self-assembly follow-up simulations for IVLFIF and VFFIF using GROMACS with AMBER99SB-ILDN and TIP3P water. Each system comprised 300 peptide copies in a $15\times 15\times 15~\mathrm{nm}^{3}$ cubic periodic box. The boxes were solvated and neutralized to electroneutrality with K+/Cl-. Energy minimization employed steepest descent (50,000 steps) with a Verlet cutoff scheme, Particle–Mesh Ewald electrostatics (1.0 nm real-space cutoff), and a 1.0 nm van der Waals cutoff. Equilibration consisted of 100 ps NVT followed by 100 ps NPT with position restraints on the peptides, using the velocity-rescale thermostat (300 K) and the Parrinello–Rahman barostat (1 bar, isotropic). Production simulations were run in the NVT ensemble for $1.25~\mu\mathrm{s}$ with a 2 fs timestep ($n_{\mathrm{steps}}=6.25\times 10^{8}$), using PME electrostatics (1.0 nm) and a 1.0 nm van der Waals cutoff; coordinates were saved every 10 ps.